Circular No. 7532

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2000ev IN UGC 3500 T. Matheson, S. Jha, P. Challis and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that an image obtained on Nov. 28.35 UT by P. Groot with the F. L. Whipple Observatory 1.2-m telescope (+ 4-Shooter) yields the following precise position for the supernova 2000ev: α = 6^{h}47^{m}52^{s}.0, δ = +84^{o}10′02′′.2 (equinox 2000.0; uncertainty 0′′.3). The supernova is 4′′.6 west and 19′′.4 north of the nucleus of UGC 3500.

Corrigendum. On IAUC 7529, line 3, for Soffrago read Sozzago.

SUPERNOVA 2000ew IN NGC 3810 M. Dennefeld and J. Patris, Institut d’Astrophysique, Paris, observed

SN 2000ew with the European Southern Observatory 3.6-m telescope and EFOSC2 (range 370–930 nm, resolution 1.8 nm) on Nov. 30.35 UT. The presence of a moderate Si II absorption, a strong 790-nm absorption and well-developed Fe II lines indicates that this is a type Ia supernova about two weeks past maximum.

SUPERNOVA 2000er IN PGC 9132 Dennefeld and Patris also followed the peculiar supernova 2000er with

the same instrument during Nov. 26–30. The R magnitude faded steadily by ∼ 0.1 mag per day, but only small changes were noticed in the spectrum since the summary by A. Maury on IAUC 7528 (of observations on Nov. 25.2 UT). A small absorption at 451 nm (rest wavelength) has disappeared, while broad emission features at 638, 658 and 708 nm have developed. The broad asymmetric emission around 589 nm is also strengthening, with its sharp blue-edge absorption (at 586 nm) not seen on any other line. It should be noted that the parent galaxy shows a strong dust lane.

Corrigendum. On IAUC 7528, line 4, for 44′′ east read 27′′ east.

SUPERNOVA 2000ej IN IC 1371 A. V. Filippenko and R. Chornock, University of California, Berkeley,

report that inspection of noisy CCD spectra (range 330–1000 nm) of SN 2000ej (IAUC 7517) obtained on Nov. 29 with the Shane 3-m reflector at Lick Observatory reveals that the object is a supernova, probably of type Ia. The spectrum most closely resembles those of subluminous members of the class, like SN 1991bg and SN 1998de, perhaps 1–2 months past maximum brightness. It should be noted that the host galaxy has a Hubble type of S0, consistent with preponderance of early-type hosts among other subluminous type Ia supernovae.

2000 November 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Brian G. Marsden